OpenStudy (anonymous):

1/2log(x)-log(1000)=3

Parth (parthkohli):

Hint: \(\log_{1000} = 2\)

Parth (parthkohli):

\[ \color{Black}{\Rightarrow {1 \over 2} \log x - 2 = 3 }\]\[ \color{Black}{\Rightarrow {1 \over 2} \log x = 5 }\]\[ \color{Black}{\Rightarrow{\log x} = 10 }\]

Parth (parthkohli):

Convert \(\log_{10}x = 10\) to exponential form and you are done.

OpenStudy (anonymous):

1/2 log(x)=3-log(1000) 1/2 log(x)=0 x=0

Parth (parthkohli):

Nope, \(x = 0\) doesn't satisfy.

OpenStudy (anonymous):

but 3-log1000 = 0

Parth (parthkohli):

What? \(\log 0 = \infty\)

Parth (parthkohli):

As I said, try to see the steps right there.

OpenStudy (anonymous):

To be pedantic, log 0 is undefined.